4. Determine the each of the currents in each path, and the voltages across each resistor. E, = 24.0 V b d 241 0.10 12 R 5.0please show which equations were used for each step
Answer #1

Using KEL I , + I2 = Is -U) Using KvL in top loop -24 to. lI +20 I, - 40 I 2 -0.SI, +48=0 SI, 25. II, 40.5I,= - 24 -(2) Using